# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

### v1

@@ 1 Introduction
Given (H1), for each $k>0$, $1\leq\ell\leq d$, and any arbitrary $r_{1}\geq r_{2}\geq 1$, then letting $\lambda_{\ell}={\lambda_{\ell}}^{\mathrm{re}}+\mathrm{i}{\lambda_{\ell}}^{\mathrm{im}}$,
$|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})|\geq C\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/{r_{2}}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}.$
@@
Write $\gamma_{k\ell}(r_{1})^{2}=\lambda_{\ell}+\mu_{k}/r_{1}^{2}=A_{1}e^{\mathrm{i}\theta_{1}}$ and $\gamma_{k\ell}(r_{2})^{2}=\lambda_{\ell}+\mu_{k}/r_{2}^{2}=A_{2}e^{\mathrm{i}\theta_{2}}$. Recall that $\lambda_{\ell}\in\mathbb{C}\backslash(-\infty,0]$ and therefore, since $\mu_{k}\in[0,\infty)$, $\theta_{1},\theta_{2}\in(-\pi+\varepsilon,\pi-\varepsilon)$ for some $\varepsilon>0$. Consequently, $\cos(\theta_{j}/2)\geq C>0$ for each $j=1,2$ where $C$ is some constant. We then estimate
$|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})|$ $=\Big|A_{1}^{1/2}\cos(\theta_{1}/2)+A_{2}^{1/2}\cos(\theta_{2}/2)+\mathrm{i}\big[A_{1}^{1/2}\sin(\theta_{1}/2)+A_{2}^{1/2}\sin(\theta_{2}/2)\big]\Big|$
$\geq A_{1}^{1/2}\cos(\theta_{1}/2)+A_{2}^{1/2}\cos(\theta_{2}/2)$
$\geq C(A_{1}^{1/2}+A_{2}^{1/2})\geq CA_{2}^{1/2}=C\left[\left({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/{r_{2}}^{2}\right)^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}$ (27)
@@
Suppose that either $A>0$ or $B\neq 0$, then for all $\xi\geq 0$,
$f(\xi):=\frac{\xi}{\left[\left(A+\xi^{2}\right)^{2}+B^{2}\right]^{1/4}}\leq C$
uniformly for some constant $C$.
@@
We begin by rewriting $\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})$ as a single fraction:
@@
$\leq|r_{1}-r_{2}|\left|\frac{r_{1}+r_{2}}{r_{1}^{2}r_{2}}\right|\left|\frac{\mu_{k}}{(\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2}))}\right|.$
We next apply Lemma A.2 to bound the denominator, and apply our assumption about the proximity of $r_{1}\geq r_{2}$ to bound the coefficients via the estimates $r_{1}+r_{2}\leq 3r_{2}$ and $r_{1}^{2}r_{2}^{2}\geq r_{2}^{4}$. This yields
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|$ $\leq C|r_{1}-r_{2}|\left|\frac{\mu_{k}/r_{2}^{3}}{\left[\left({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{2}^{2}\right)^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}}\right|.$ (28)
@@
Let $r\geq 1$ and consider $r^{\prime}\geq 1$ sufficiently close to $r$ so that $r^{\prime}/2\leq r<2r^{\prime}$. Then, given (H2), for all $k>0$ and $1\leq\ell\leq d$,
$|\gamma_{k\ell}(r)-\gamma_{k\ell}(r^{\prime})-\partial_{r}\gamma_{k\ell}(r^{\prime})(r-r^{\prime})|\leq C(1+\mu_{k})^{1/2}|r-r^{\prime}|^{2}$
where $\partial_{r}\gamma_{k\ell}$ is the $r$ derivative of $\gamma_{k\ell}$ and also the symbol for the Fréchet derivative of $A(r)^{1/2}$.
@@
The result follows from a rather tedious calculation in which we suppress the $\ell$ subscript:
$\gamma_{k}(r)$- $\gamma_{k}(r^{\prime})-\partial_{r}\gamma_{k}(r^{\prime})(r-r^{\prime})=\frac{\mu_{k}(1/r^{2}-1/r^{\prime 2})}{\gamma_{k}(r)+\gamma_{k}(r^{\prime})}-\frac{\mu_{k}(r^{\prime}-r)}{(r^{\prime})^{3}\gamma_{k}(r^{\prime})}$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{r^{\prime}+r}{r^{2}(\gamma_{k}(r)+\gamma_{k}(r^{\prime}))}-\frac{1}{r^{\prime}\gamma_{k}(r^{\prime})}\right]$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{r^{\prime}(r+r^{\prime})\gamma_{k}(r^{\prime})-r^{2}(\gamma_{k}(r)+\gamma_{k}(r^{\prime}))}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right]$
$=(r^{\prime}-r)\frac{\mu_{k}}{(r^{\prime})^{2}}\left[\frac{(r^{\prime})^{2}(\gamma_{k}(r^{\prime})-\gamma_{k}(r))+\gamma_{k}(r)((r^{\prime})^{2}-r^{2})+\gamma_{k}(r^{\prime})r(r^{\prime}-r)}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right]\quad\bigg(\begin{matrix}\text{add and subtract}\\ (r^{\prime})^{2}\gamma_{k}(r)\end{matrix}\bigg)$
$=-(r^{\prime}-r)^{2}\frac{\mu_{k}}{(r^{\prime})^{2}}\Bigg[\frac{(r+r^{\prime})\mu_{k}}{r^{\prime}r^{4}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}-\frac{\gamma_{k}(r)(r^{\prime}+r)}{r^{\prime}r^{2}\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}$
$\qquad\qquad\qquad\qquad\quad-\frac{1}{r^{\prime}r(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\Bigg]$
$=-(r-r^{\prime})^{2}\frac{\mu_{k}^{1/2}}{r^{\prime}}\Bigg[\frac{r^{\prime}(r+r^{\prime})}{r^{4}}\frac{\mu_{k}^{3/2}/(r^{\prime})^{3}}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}-\frac{(r+r^{\prime})}{r^{\prime}r^{2}}\frac{\gamma_{k}(r)(\mu_{k}^{1/2}/r^{\prime})}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}$
$\qquad\qquad\qquad\qquad\quad-\frac{1}{r^{\prime}r}\frac{\mu_{k}^{1/2}/r^{\prime}}{\gamma_{k}(r^{\prime})+\gamma_{k}(r)}\Bigg]$
[…]
We can then bound $|I|$, $|II|$, and $|III|$ using the same approach as in the proof of Lemma A.4. In particular, using Lemma A.2, Lemma A.3, and the relationship between $r$ and $r^{\prime}$ we obtain:
$|I|$ $\leq 3\cdot 2^{4}\left|\frac{\mu_{k}^{3/2}/r^{\prime 2}}{\gamma_{k}(r^{\prime})(\gamma_{k}(r^{\prime})+\gamma_{k}(r))^{2}}\right|\leq C$
$|II|$ $\leq 3\cdot 4\left|\frac{\gamma_{k}(r)}{\gamma_{k}(r^{\prime})}\right|\left|\frac{\mu_{k}^{1/2}/r^{\prime}}{(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right|\leq C$
$|III|$ $\leq 2\left|\frac{\mu_{k}^{1/2}/r^{\prime}}{(\gamma_{k}(r^{\prime})+\gamma_{k}(r))}\right|\leq C$
for some constant $C>0$ uniformly in $k$. ∎
@@
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|\leq C(1+\mu_{k})^{1/2}\left(|r_{1}-r_{2}|+|r_{1}^{2}-r_{2}^{3}|\right).$
@@
We begin by writing the difference as a single fraction and simplifying.
@@
$=\mu_{k}\left|\frac{r_{1}^{3}(\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2}))-\gamma_{k\ell}(r_{2})(r_{2}^{3}-r_{1}^{3})}{r_{1}^{3}r_{2}^{3}\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})}\right|\qquad\text{(add and subtract $r_{1}^{3}\gamma_{k\ell}(r_{2})$ )}$
$\leq\left(\frac{\mu_{k}}{r_{2}^{3}|\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})|}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}}{r_{1}^{3}r_{2}^{3}|\gamma_{k\ell}(r_{1})|}\right)\left|r_{2}^{3}-r_{1}^{3}\right|$
Applying Lemma A.2 for $r_{1}=r_{2}$, we obtain
$2|\gamma_{k\ell}(r_{j})|\geq\tilde{C}\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{j}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}$
$\text{and}\qquad\frac{\mu_{k}^{1/2}/r_{j}}{|\gamma_{k\ell}(r_{j})|}\leq 2\cdot\frac{\mu_{k}^{1/2}/r_{j}}{\tilde{C}\left[({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r_{j}^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}\right]^{1/4}}$ $\leq C$
for each $j=1,2$. Furthermore, by our assumption that $r_{1}$ and $r_{2}$ are sufficiently close, we have
$\frac{\mu_{k}^{1/2}/r_{2}}{\gamma_{k\ell}(r_{1})}\leq\frac{2\mu_{k}^{1/2}/r_{1}}{\gamma_{k\ell}(r_{1})}\leq 2C.$
@@
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $\leq C\left[\left(\frac{\mu_{k}^{1/2}}{r_{2}^{2}|\gamma_{k\ell}(r_{1})|}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}^{1/2}}{r_{1}^{2}r_{2}^{3}}\right)\left|r_{1}^{3}-r_{2}^{3}\right|\right]$
$\leq C\left[\left(\frac{2C}{r_{2}}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}^{1/2}}{r_{1}^{2}r_{2}^{3}}\right)\left|r_{1}^{3}-r_{2}^{3}\right|\right]$
$\leq C\left[2C\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\mu_{k}^{1/2}\left|r_{1}^{3}-r_{2}^{3}\right|\right],\qquad\quad(r_{1},r_{2}\geq 1).$
@@
For any $u\in H^{1}(S^{n-1},\mathbb{C})$, there is a constant $C>0$ and independent of $r$ such that
$\frac{1}{r^{2}}\|u\|_{H^{1}}^{2}\leq C\|u\|_{L^{2}}^{2}+\|A(r)^{1/2}u\|_{L^{2}}^{2}.$
Furthermore, if (H2) holds, then for some constant $C$ independent of $r$,

### v2

@@ 1 Introduction
Given (H2), for each $k\geq 0$, $1\leq\ell\leq d$, and any arbitrary $r_{1},r_{2}\geq 1$,
$|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})|\geq|\gamma_{k\ell}(r_{j})|,\qquad j=1,2.$
@@
Recall that $\gamma_{k\ell}(r)^{2}=\lambda_{\ell}+\mu_{k}/r^{2}$ where $\mu_{k}\geq 0$ and $\lambda_{\ell}\in\mathbb{C}\backslash(-\infty,0]$. If $\mathrm{Im}\left[\lambda_{\ell}\right]\geq 0$, then by taking the principal branch of the square root, we also have $\mathrm{Im}\left[\gamma_{k\ell}(r)\right]\geq 0$ for all $r\geq 1$. Similarly, if $\mathrm{Im}\left[\lambda_{\ell}\right]\leq 0$, then $\mathrm{Im}\left[\gamma_{k\ell}(r)\right]\leq 0$ for all $r\geq 1$. Since for each $\ell$ the imaginary parts of $\gamma_{k\ell}(r_{1})$ and $\gamma_{k\ell}(r_{2})$ have matching sign, we have that
$\mathrm{Im}\left[\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})\right]^{2}=\left(\mathrm{Im}\left[\gamma_{k\ell}(r_{1})\right]+\mathrm{Im}\left[\gamma_{k\ell}(r_{2})\right]\right)^{2}\geq\mathrm{Im}\left[\gamma_{k\ell}(r_{j})\right]^{2}.$
Furthermore, because we are taking the principal branch of the square root, $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]>0$ for all $r$ and all $\ell$. Therefore we have
$\mathrm{Re}\left[\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})\right]^{2}=\left(\mathrm{Re}\left[\gamma_{k\ell}(r_{1})\right]+\mathrm{Im}\left[\gamma_{k\ell}(r_{2})\right]\right)^{2}\geq\mathrm{Re}\left[\gamma_{k\ell}(r_{j})\right]^{2}.$
Consequently, we have that
$\left|\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})\right|$ $=\sqrt{\mathrm{Re}\left[\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})\right]^{2}+\mathrm{Im}\left[\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})\right]^{2}}$
$\geq\sqrt{\mathrm{Re}\left[\gamma_{k\ell}(r_{j})\right]^{2}+\mathrm{Im}\left[\gamma_{k\ell}(r_{j})\right]^{2}}=|\gamma_{k\ell}(r_{j})|.$
@@
Given (H2), for $k>0$, $1\leq\ell\leq d$, and any $r_{1},r_{2}\geq 1$,
$\left|\frac{\mu_{k}^{1/2}/r_{j}}{\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})}\right|\leq C,\qquad j=1,2$
where $C>0$ is some constant independent of $k$, $\ell$, and $r_{1,2}$.
Proof.
By Lemma A.2, we have that
$\left|\frac{\mu_{k}^{1/2}/r_{j}}{\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})}\right|^{2}\leq\frac{\mu_{k}/r_{j}^{2}}{|\gamma_{k\ell}(r_{j})|^{2}}=\frac{\mu_{k}/r_{j}^{2}}{\big|\gamma_{k\ell}(r_{j})^{2}\big|}=\frac{\mu_{k}/r_{j}^{2}}{\big|\lambda_{\ell}+\mu_{k}/r_{j}^{2}\big|}=f_{\ell}\left(\mu_{k}/r_{j}^{2}\right).$
where $f_{\ell}(\xi):=\dfrac{\xi}{|\lambda_{\ell}+\xi|}$.
Notice that $f_{\ell}(\xi)$ is well-defined and continuous on the domain $\xi\in[1,\infty)$ since $\lambda_{\ell}\in\mathbb{C}\backslash(-\infty,0]$, and that $\lim_{\xi\to\infty}f(\xi)=1$. Furthermore, it is straightforward to show using calculus techniques that $f_{\ell}(\xi)$ is strictly increasing for $\xi\geq 1$ provided $\mathrm{Re}\left[\lambda_{\ell}\right]\geq 0$. If $\mathrm{Re}\left[\lambda_{\ell}\right]<0$, then by solving $f^{\prime}_{\ell}(\xi)=0$ we find that $f_{\ell}(\xi)$ attains a single maximum on $1,\infty)$ at $\xi=-|\lambda_{\ell}|^{2}/\mathrm{Re}\left[\lambda_{\ell}\right]$ given by
$f_{\ell}\left(\frac{|\lambda_{\ell}|^{2}}{\left|\mathrm{Re}\left[\lambda_{\ell}\right]\right|}\right)=\frac{|\lambda_{\ell}|}{\left|\mathrm{Im}\left[\lambda_{\ell}\right]\right|}.$
The result follows with constant $C>0$ chosen as
$C=\max_{1\leq\ell\leq d}\left(\sup_{\xi\geq 1}\sqrt{f_{\ell}(\xi)}\right).$
@@
We begin by multiplying $\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})$ by $(\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2}))$ in both the numerator and the denominator, and simplifying the result into a single fraction:
@@
$\leq|r_{1}-r_{2}|\left|\frac{r_{1}+r_{2}}{r_{1}^{2}r_{2}^{2}}\right|\left|\frac{\mu_{k}}{(\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2}))}\right|.$
We next apply our assumption that $r_{2}\leq r_{1}\leq 2r_{2}$ to bound the coefficients via the estimates $r_{1}+r_{2}\leq 3r_{2}$ and $r_{1}^{2}r_{2}^{2}\geq r_{2}^{4}$. For $C=3$, this yields
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|$ $\leq C|r_{1}-r_{2}|\left|\frac{\mu_{k}/r_{2}^{3}}{\gamma_{k\ell}(r_{1})+\gamma_{k\ell}(r_{2})}\right|.$
@@
Let $r_{1},r_{2}>0$ be sufficiently close so that $r_{1}/2\leq r_{2}<2r_{1}$. Then, given (H2), for all $k>0$ and $1\leq\ell\leq d$,
$|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})-\partial_{r}\gamma_{k\ell}(r_{1})(r_{1}-r_{2})|\leq C(1+\mu_{k})^{1/2}|r_{1}-r_{2}|^{2}.$
@@
The result follows from a rather tedious calculation in which we suppress the $\ell$ subscript. Notice that in the first step of the calculation we simplify $(\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2}))$ as in the proof of Lemma A.4.
$\gamma_{k}(r_{1})$- $\gamma_{k}(r_{2})-\partial_{r}\gamma_{k}(r_{1})(r_{1}-r_{2})=\frac{\mu_{k}(r_{2}+r_{1})(r_{2}-r_{1})}{r_{1}^{2}r_{2}^{2}(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}-\frac{\mu_{k}(r_{2}-r_{1})}{r_{1}^{3}\gamma_{k}(r_{1})}$
$=(r_{2}-r_{1})\frac{\mu_{k}}{r_{1}^{2}}\left[\frac{r_{2}+r_{1}}{r_{2}^{2}(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}-\frac{1}{r_{1}\gamma_{k}(r_{1})}\right]$
$=(r_{2}-r_{1})\frac{\mu_{k}}{r_{1}^{2}}\left[\frac{r_{1}(r_{1}+r_{2})\gamma_{k}(r_{1})-r_{2}^{2}(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}{r_{1}r_{2}^{2}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right]$
$=(r_{2}-r_{1})\frac{\mu_{k}}{r_{1}^{2}}\left[\frac{r_{1}^{2}(\gamma_{k}(r_{1})-\gamma_{k}(r_{2}))+\gamma_{k}(r_{2})(r_{1}^{2}-r_{2}^{2})+r_{2}\gamma_{k}(r_{1})(r_{1}-r_{2})}{r_{1}r_{2}^{2}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right]$
We pause to note that this last equality is obtained by adding and subtracting the term $r^{2}_{1}\gamma_{k}(r_{2})$. Continuing with the calculation, this last term is equivalent to:
$(r_{2}-r_{1})$ $\frac{\mu_{k}}{r_{1}^{2}}\Bigg[\frac{\mu_{k}(r_{2}+r_{1})(r_{2}-r_{1})}{r_{1}r_{2}^{4}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))^{2}}+(r_{1}-r_{2})\left(\frac{\gamma_{k}(r_{2})(r_{1}+r_{2})+r_{2}\gamma_{k}(r_{1})}{r_{1}r_{2}^{2}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right)\Bigg]$
$=$ $(r_{2}-r_{1})^{2}\frac{\mu_{k}}{r_{1}^{2}}\Bigg[\frac{\mu_{k}(r_{2}+r_{1})}{r_{1}r_{2}^{4}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))^{2}}$
$\hskip 79.49744pt-\frac{\gamma_{k}(r_{2})(r_{1}+r_{2})}{r_{1}r_{2}^{2}\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}-\frac{1}{r_{1}r_{2}(\gamma_{k}(r)+\gamma_{k}(r_{2}))}\Bigg]$
$=$ $(r_{2}-r_{1})^{2}\frac{\mu_{k}^{1/2}}{r_{1}}\Bigg[\frac{r_{1}(r_{2}+r_{1})}{r_{2}^{4}}\frac{\mu_{k}^{3/2}/r_{1}^{3}}{\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))^{2}}$
$\hskip 79.49744pt-\frac{(r_{2}+r_{1})}{r_{1}r_{2}^{2}}\frac{\gamma_{k}(r_{2})(\mu_{k}^{1/2}/r_{1})}{\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}-\frac{1}{r_{1}r_{2}}\frac{\mu_{k}^{1/2}/r_{1}}{(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\Bigg]$
From this calculation we have that
$\big|\gamma_{k}(r_{1})$- $\gamma_{k}(r_{2})-\partial_{r}\gamma_{k}(r_{1})(r_{1}-r_{2})\big|\leq(r_{1}-r_{2})^{2}\frac{\mu_{k}^{1/2}}{r_{1}}\Big[|I|+|II|+|III|\Big]$
where, by Lemma A.3, the assumption that $r_{1}/2\leq r_{2}\leq 2r_{1}$, and the continuity of $\gamma_{k}$ on $r_{1}\geq 1$, we have
$|I|=\left(\frac{r_{1}(r_{2}+r_{1})}{r_{2}^{4}}\right)\left|\frac{\left(\mu_{k}^{1/2}/r_{1}\right)^{3}}{\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))^{2}}\right|$ $\leq 3\cdot 2^{4}\,\left|\frac{\left(\mu_{k}^{1/2}/r_{1}\right)^{3}}{\gamma_{k}(r_{1})(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))^{2}}\right|\leq C$
$|II|=\left(\frac{r_{2}+r_{1}}{r_{1}r_{2}^{2}}\right)\left|\frac{\gamma_{k}(r_{2})}{\gamma_{k}(r_{1})}\right|\left|\frac{\mu_{k}^{1/2}/r_{1}}{(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right|$ $\leq 3\cdot 4\left|\frac{\gamma_{k}(r_{2})}{\gamma_{k}(r_{1})}\right|\left|\frac{\mu_{k}^{1/2}/r_{1}}{(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right|\leq C$
$|III|=\quad\left(\frac{1}{r_{1}r_{2}}\right)\quad\left|\frac{\mu_{k}^{1/2}/r_{1}}{(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right|\quad$   $\leq\,\,2\,\,\left|\frac{\mu_{k}^{1/2}/r_{1}}{(\gamma_{k}(r_{1})+\gamma_{k}(r_{2}))}\right|\leq C.$
Finally, the simple bound in Lemma A.1, namely, $\mu_{k}^{1/2}/r_{1}\leq(1+\mu_{k})^{1/2}$, concludes the proof. ∎
@@
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|\leq C(1+\mu_{k})^{1/2}\left(|r_{1}-r_{2}|+|r_{1}^{3}-r_{2}^{3}|\right).$
@@
We begin by writing the difference as a single fraction.
@@
$=\mu_{k}\left|\frac{r_{1}^{3}(\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2}))-\gamma_{k\ell}(r_{2})(r_{2}^{3}-r_{1}^{3})}{r_{1}^{3}r_{2}^{3}\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})}\right|$
The second line of the calculation above is obtained by adding and subtracting $r_{1}^{3}\gamma_{k\ell}(r_{2})$ in the numerator. Breaking up this fraction and applying triangle inequality we obtain:
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $\leq\left(\frac{\mu_{k}}{r_{2}^{3}|\gamma_{k\ell}(r_{2})\gamma_{k\ell}(r_{1})|}\right)\!\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\!\left(\frac{\mu_{k}}{r_{1}^{3}r_{2}^{3}|\gamma_{k\ell}(r_{1})|}\right)\left|r_{2}^{3}-r_{1}^{3}\right|.$
Applying the same argument as in Lemma A.3 we obtain
$\frac{\mu_{k}^{1/2}/r_{j}}{|\gamma_{k\ell}(r_{j})|}\leq C,\qquad j=1,2.$
Furthermore, by our assumption that $r_{2}/2\leq r_{1}\leq 2r_{2}$, we have
$\frac{\mu_{k}/r_{2}^{2}}{|\gamma_{k\ell}(r_{1})\gamma_{k\ell}(r_{2})|}\leq\left(\frac{2\mu_{k}^{1/2}/r_{1}}{|\gamma_{k\ell}(r_{1})|}\right)\left(\frac{\mu_{k}^{1/2}/r_{2}}{|\gamma_{k\ell}(r_{2})|}\right)\leq C.$
@@
$|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|$ $\leq C\left[\left(\frac{1}{r_{2}}\right)\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\left(\frac{\mu_{k}^{1/2}}{r_{1}^{2}r_{2}^{3}}\right)\left|r_{1}^{3}-r_{2}^{3}\right|\right]$
$\leq C\left[\left|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})\right|+\mu_{k}^{1/2}\left|r_{1}^{3}-r_{2}^{3}\right|\right],\qquad\quad(r_{1},r_{2}\geq 1).$
@@
For any $u\in H^{1}(S^{n-1},\mathbb{C})$, there is a constant $C\geq 0$ and independent of $r$ such that
$\frac{1}{r^{2}}\|u\|_{H^{1}}^{2}\leq 2C\|u\|_{L^{2}}^{2}+2\|A(r)^{1/2}u\|_{L^{2}}^{2}.$
Furthermore, if both (H2) and (H3) hold, then for some constant $C>0$ independent of $r$,